OpenStudy (anonymous):

What is the equation of a circle with a center at (-4, -8) and a radius of 1.9 ?

Vocaloid (vocaloid):

(x-h)^2 + (y-k)^2 = r^2 is the standard equation of a circle, where h and k are the x and y coordinates, respectively, and r is the radius. Substituting the given values for r, h, and k will give you the equation of the circle in standard form.

OpenStudy (anonymous):

@Vocaloid (x+40^2 + (y+8)^2=1.9^2 would be the final equation corect?

Vocaloid (vocaloid):

Yes (I assume the 0 is a typo for a parentheses?)

OpenStudy (anonymous):

yes it was

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Vocaloid